OpenStudy (anonymous):

The cost per minute of a conference phone call, originally $2.80, increased by 20%. What is the new price? A. $3.00 B. $3.20 C. $3.26 D. $3.36

OpenStudy (anonymous):

okay can some one help me other then say that is wrong thx boys

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2.80 * 120% = 3.36

OpenStudy (anonymous):

so D.

OpenStudy (anonymous):

U got it?

OpenStudy (nubeer):

how much price increased = 20% of 2.8 = 0.2*2.8 =0.56 new price = 2.8+(0.2*2.8)

OpenStudy (anonymous):

same thing..

OpenStudy (nubeer):

i know.. never complained with your method.

OpenStudy (anonymous):

thx